buffer chosen: phosphate, pH before addition of base 7.21 after addition of base 8.87

1) What is the ratio of A-/HA in your buffer after adding acid or base?

2) How many micromoles of A- & HA are present in the solution?

3) Calculate the theoretical pH of the buffer after adding .5mL of 1M HCl or 1M NaOH.
Answer #1

1) According to Henderson-Hasselbulch equation:

pH = pKa + Log([A-]/[HA])

i.e. 7.21 = 7.2 + Log([A-]/[HA])

i.e. Log([A-]/[HA]) = 7.21 - 7.2 = 0.01

i.e. [A-]/[HA] = 100.01 = 1.023 = 1.023/1

2) No. of micromoles of A- and HA present in the solution = 1023 + 1000 = 2023 micromoles

3) a) Addition of strong acid = 0.5 mL * 1 mmol/mL = 0.5 mmol

pH = pKa + Log{([A-] - [Strong acid]/([HA] + [strong acid])}

i.e. pH = 7.2 + Log{(1.023 - 0.5)/(1 + 0.5)}

i.e. pH = 6.74

b) Addition of strong base = 0.5 mL * 1 mmol/mL = 0.5 mmol

pH = pKa + Log{([A-] + [Strong base]/([HA] - [strong base])}

i.e. pH = 7.2 + Log{(1.023 + 0.5)/(1 - 0.5)}

i.e. pH = 7.68